The lenght of a rectangle is (3x) yards and its width is (x+2) yards. For what value of x will the perimeter of the rectangle be 40 yards.
Answer by calbreez(9)

Perimeter = 2(length+breadth)
=2(3x+x+2)
=2(4x+2)
=8x+4
Since the perimeter is required to be 40yds,
8x+4=40
8x=36
x=4.5